Sol Summary – May 22nd

Crew 281 Sol Summary Report 22 May 2023
Sol: 8
Summary Title: Hair wash day
Author’s name: Ritupriya Patil (XO) and Megan Kane (Commander)
Mission Status: Nominal
Sol Activity Summary:
Today’s EVA was at Pooh’s corner. The crew of 3 drove Spirit and Perseverance rovers and reached the site at 0823 hrs. They captured scanline measurements for the boulders Ana Pires (Crew Scientist) approved at Pooh’s Corner. KC Shasteen helped with the Equotip measurements while Ana gathered the Schmidt hammer data. The crew finished a 5 station long scanline and during their measurements Ritupriya Patil (XO, EVA Lead) deployed the drone for aerial imagery. Detailed vertical images and videos of the scanline were captured to help confirm the weathering grade, discontinuities of the rocks. These images and videos will be used for georeferencing the area around the scanline and for further assessment of the geological region. The objective was achieved right before winds gusted to 27 kmph and the drone was secured.
For lunch Ritu (XO) made tofu Thai curry with white rice. Post lunchtime Ritu and Ana worked on tomorrow’s EVA planning, looking at maps, distance and noting time for each stop over during the 2 person EVA. Then Ritu and KC worked in the RAM to build a drone packaging so it can be securely stored during the travel time in a ready to be deployed configuration. This makes the drone launch much faster and efficient.
Rachel (HSO) checked into OMISS Old Man International Side band Society through the ham radio setup. She is working through troubleshooting the power reception she is having. While Ana worked on closing the set of scanlines and working further on research and synergies with Ritu. KC and Rachel had a hair wash day to clear their heads (get that?!).
In the green hab, Megan (Commander) watered and checked on the plants. KC (Crew Engineer) did routine checks on hab operations – batteries, tunnels, flushed toilet and water levels. Ana cleaned up her robotic scoop in the RAM after a successful mission today.
In the science dome Megan checked on the cacao seeds and made a video recording the changes to the seeds.
For dinner Rachel made stamppotten a mashed potato dish from the Netherland. It rained in the late afternoon and there were storms, but the antenna stayed upright.
Look Ahead Plan: Tomorrow we will do a 2-person, long range EVA to the Overlook region. It will be an all-female crew exploring the Martian surface and capturing aerial footage. On our way back we will stop at Pooh’s Corner for images of scanlines that were measured on earlier EVAs.
Anomalies in work: None
Weather: Windy
Crew Physical Status: Nominal
EVA: Completed 1 today

Reports to be filed:
Greenhab Report
EVA Report
EVA Request
Ops Report
Journalist Report
Photo of the Day

Support Requested:
Mission support is requested in case of storms if the antenna needs to come down or be retrieved.

Operations Report – May 22nd

Crew 281 Operations Report 22 May 2023
SOL: 8
Name of person filing report: Megan Kane (Commander), Ritupriya Patil (Executive Officer) and KC Shasteen (Crew Engineer)
Non-nominal systems:
Electrical outlets in kitchen
Remote observatory
EVA Helmet
Notes on non-nominal systems: None
Spirit rover used: Yes
Hours: (before EVA): 222.2
Hours: (after EVA): 222.5
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 78
Currently charging: Yes
Opportunity rover used: No
Hours: (before EVA): 130.0
Hours: (after EVA): NA
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): NA
Currently charging: Yes
Curiosity rover used: No
Hours: (before EVA): 240.2
Hours: (after EVA): NA
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): NA
Currently charging: Yes
Perseverance rover used: Yes
Hours: (before EVA): 264.3
Hours: (after EVA): 264.6
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 66%
Currently charging: Yes
General notes on rovers: None
Summary of Hab operations:
WATER USE: 39.9 gallons
Water (static tank): 281.9 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: yes
Summary of internet:
Summary of suits and radios: All suits and radios are fully recharged.
Summary of GreenHab operations:
WATER USE: 2.64 gallons
Heater: Off
Supplemental light: Off
Harvest: in grams : 0
Summary of Science Dome operations:
Dual split: Heat and AC, Off
EVA & Research Planning
Cacao Experiment:
Watered cacao seedlings
Summary of RAM operations: (Tools used, work done)
Fixed Patches for Ana and KC
Fabricated restraint for carrying drone on EVA
Summary of any observatory issues: NA
Summary of health and safety issues:
No new issues to report.
Questions, concerns and requests to Mission Support:
Request Mission Support assistance to take down or retrieve the ham radio antenna in case of emergency or high winds.

Journalist Report – May 22nd

Crew 281 Journalist Report 22 May 2023

Journalist Report

By: Rachel Jones

Sol 8, Getting Our Procedures Straight.

It’s Sol 8, and things are starting to feel like clockwork. Everyone knows I’ll ask them if they have water, if they’ve been drinking water, if they brought extra water, etc. etc. etc. Our EVA of Ana, Ritu, and KC went to Pooh’s Corner test to perform the last scan line and take aerial footage. Our Crew has taken over 950 measurements of rock hardness with the Equotip and Schmidt Hammer. The crew has the system down to a point where neither the slightly damp ground nor the poor band signal interfered with the mission.

Lunch today, prepared by Ritu, was Tofu Thai Curry and white rice. I spent my lunchtime hour trying to check into the Old Man International Sideband Society’s (OMISS) daily net. For those not into amateur radio, a “net” is traditionally an on-the-air gathering. I was on 20 meters and could barely be heard. I plan to try again tomorrow and see if I can get through. You can check them out on

After lunch, the great hair-washing experiment commenced. Our goal was to use as little water as possible. We averaged about 3/4s of a gallon per person. We then reused the hair-washing water to top off and flush the toilet system. From cooking to basic chores, we like to make every drop count. After my hair was washed, Megan our impromptu stylist, put it up in a French braid.

The afternoon sent everyone off to their various tasks. I collaborated with various radio operators to try and diagnose my radio set up to be clearer. KC and Ritu constructed a new on-EVA drone transportation system so that the drone could be carried in a ready-to-be-deployed configuration (gloves make it hard to do in the field). Ana and Ritu then planned out tomorrow’s EVA. Megan took care of the GreenHab and her chocolate plants. The chocolate plants are starting to sprout now.

Dinner tonight was up to me. For all those of you that have waited with bated breath to learn what-all is involved in cooking at MDRS, wait no longer. Given the variety of options we had, I chose to make a MDRS version of stamppotten, a potato and vegetable dish. First, I gathered the various vegetables I wanted to include. Each ingredient had a different amount of time it needed to rehydrate. I put about two tablespoons of the first two ingredients (carrot flakes and cheese gratings) in a large bowl and added warm water. They needed 15 mins to rehydrate. Five minutes later, I added the next two veggies (onion flakes and peppers bits). Then 5 minutes later I added the last two ingredients (spinach flakes and butter powder). On the stove, I put in
2 cups of the instant mashed potatoes mix, salt, garlic, and some of the rehydrated milk mix. Then I added the hydrated vegetables (along with the excess water) into the potatoes and brought it to a quick boil. I think it turned out fantastic.

Feeling full and happy, I bid you all ado for tonight.

GreenHab Report – May 22nd

Crew 281 GreenHab Report 22 May 2023
GreenHab Officer: Megan Kane
Environmental control: mixed door open/ door open with fan
Average temperatures: 87.78 °F
Hours of supplemental light: None
Daily water usage for crops 2.25 gallons
Daily water usage for research and/or other purposes: 0.4 gallons
Water in Blue Tank 161.12 gallons
Time(s) of watering for crops:
0720: refilled 5 water bottle (2.5 L) watered other plants as needed 3.0 L
1813: refill 3 water bottles (1.5L) and watered other plants as needed 1.5L
Changes to crops:
Tomatoes and cucumbers almost ready to harvest.
In the morning the GreenHab was checked and watered. The cooling cloth’s reservoir was also refilled. The temperature was checked several times. The highest recorded temperature today was 102.7.
Harvest: (include which crop and mass in grams): None
Support/supplies needed: None

EVA Report – May 22nd

Crew 281 EVA Report 22-05-2023
EVA # 9
Authors: Megan Kane (Commander), Ritupriya Patil (Executive Officer), Ana Pires (Crew Scientist)
Purpose of EVA:
Conduct Geotechnical Survey of the area and geomechanical assessment
Schmidt Hammer and Equotip measurements along scanlines
Geological and geotechnical assessment of the scanline using the created datasheet by the Crew Scientist
Acquire images and collect loose rock/soil samples
Drone flight to record all of the above operations and do aerial imagery
Start time: 0834 hrs
End time: 0939 hrs
At 0801 the three members of the EVA entered the airlock exiting at 0807. They first checked all the rovers and collected their information. Hours and State of Charge were recorded for each vehicle and it was verified that all were securely plugged in. They then unplugged Spirit and Perseverance in preparation for heading out to Pooh’s Corner (Site). They left at 0814.
EVA Crew arrived at Site 0823. After parking the rovers the crew did a visual survey of the area to locate a good location for the scanline. At 0806 the crew conducted a water break. They chose the location and laid the scanline. Measurements commenced at approximately 0834. There were 5 stations total. See the completion times below:
Station 1 complete at 0838.
Station 2 complete at 0851
Station 3 complete at 0858
Station 4 complete at 0906
Station 5 complete at 0910
Water checks were taken approximately every half an hour. Samples were collected near the scanline after station 5 was complete. When station 3 began at 0840 drone setup was started. Drone was launched at 0840. The wind was very high and the drone was unable to complete the planned shots. The drone landed at 0900.
At 0919 the crew headed back to the rovers. At 0922 they started back to MDRS in rovers arriving at 0930. The crew re-entered the airlock at 0934 and the hab at 0939.
The crew did not stop over at Marble Ritual for drone footage due to high winds.
Overview of EVA
Drove to Pooh’s Corner
Completed a scanline
Took drone footage for georeferencing

Issues Experienced During EVA
Comms issues throughout. We expect this is due to the high inference on the bands from the current space weather.
Winds too high to get all desired imagery.

Crew completed Rover checks
Crew drove to Pooh’s Corner Ridge
Crew Completed Geological survey
Drone footage of survey area at Pooh’s Corner collected
Collected 6 rocks, 936 g; 1 tube of soil, 49 g

Coordinates (use UTM WSG 84): 519025 E, 4251183 N
EVA Participants 3: Ritupriya Patil (XO, EVA Lead, KC Shasteen (Crew Engineer), Ana Pires (Crew Scientist)
Road(s) and routes per MDRS Map: Drive on Cow Dung Road and walk east until Pooh’s Corner.
Mode of travel: Rover

Copyright © The Mars Society. All rights reserved. | Main Site